(-3/y+5)-(4/y-5)=2/y^2-25
Answer by stanbon(75887)

(-3/y+5)-(4/y-5)=2/y^2-25
Multiply thru by y^2-25 to get:
-3(y-5)-4(y+4)=2
-7y-1=2
7y=-3
y=-3/7
Cheers,
Stan H.